# Involutive (simple) latin solutions of the Yang-Baxter equation and related (left) quasigroups

### v1

@@ Introduction
The seminal papers due by Gateva-Ivanova and Van Den Bergh [25] and Etingov, Schedler, and Soloviev [24] encouraged several authors to the study of the involutive non-degenerate set-theoretic solutions (which we simply call involutive solutions). Recall that a solution $(X,r)$ is said to be decomposable if $X$ can be expressed as a disjoint union of two non-empty subsets $Y_{1}$ and $Y_{2}$ such that $r(Y_{i}\times Y_{i})\subseteq Y_{i}\times Y_{i}$ for all $i\in\{1,2\}$, and $(X,r)$ is called indecomposable, otherwise. To attack the classification-problem of involutive solutions, as first attempt one can restrict the investigations to the class of indecomposable ones. This approach is motivated by the fact that indecomposable solutions are in some sense the fundamental blocks that allow to understand all the involutive ones (see [24, Section $3$]). Even if we focus on indecomposable involutive solutions, a full classification is still hard, and in the last few years several authors investigated the so-called simple indecomposable involutive solutions, where a solution $(X,r)$ is said to be simple if every epimorphic image $(Y,s)$ is isomorphic to $(X,r)$ or is a singleton. A notion of simple involutive solution (not necessarily indecomposable) was at first introduced by Vendramin in [41]. Even if his definition of simplicity is different, it coincides in the indecomposable case. Every simple involutive solution of size bigger than $2$ is indecomposable by [16]. These solutions play a special role since they are the ”fundamental blocks” to construct all the indecomposable involutive ones by dynamical extensions (see [11, Proposition 2] and [41, Corollary 2.13]).
Even if the construction of simple involutive solutions appeared quite hard, several interesting results were obtained in the last few years. If $p$ is an odd prime number there exist a unique simple involutive solution of size $p$, see [24, Theorem 2.13]. For every natural number $n>1$, a simple involutive solution of cardinality $p^{n}$ was recently constructed in [18], and in the case $n=2$ a full classification was given in [21]. Further examples of cardinality $n^{2}$ and $mn^{2}$, where $n,m$ are arbitrary natural number, were constructed in [16, 19]. On the other hand we know that an indecomposable involutive solutions of square-free cardinality can not be simple as showed in [17].
@@
holds for all $x,y,z\in B$. The notion of simplicity was recently investigated in the general setting of bijective non-degenerate solutions (simply called solution) not necessarily involutive. In [12], a criterion of simplicity was provided by blocks of imprimitivity of the associated permutation group. A remarkable generalization of [9, Theorem $12$] was recently provided in [20], where the simple bijective solutions (not necessarily involutive) were completely characterised by the underlying permutation skew left braces, and several examples of non-involutive simple solutions were provided. In this way, the main Theorem of [20, Section $1$] represents a unifying tool to understand the notion of simplicity of several algebraic structures already present in literature. By these last results, it is clear that for the simplicity of a solution the ideal $B^{(2)}$ of the permutation left brace plays a fundamental role. However, we are still far away from a concrete description of the maps $r$ related to simple solutions. Indeed, in [20] a complete description of simple solutions provided by skew left braces $B$ with $B^{(2)}=0$ or $B^{(3)}=0$ was given, but the remaining case $B^{(2)}=B^{(3)}\neq 0$ is still open. As noted in [20, Section $1$], the last case includes the involutive simple solutions that are not of Lyuabashenko type. To partially fill this gap, one of the main goals of this paper is the study of finite involutive simple solutions in which $B^{(2)}$, that in this case coincides with the so-called displacement group, acts regularly on the underlying set. We will show that these solutions belong to the family of latin solutions, i.e. solutions in which the maps $\lambda_{x}$ provide a quasigroup structure on the set $X$ (see [8]).However, our investigation is not limited to simple solutions. Motivated by [15, Theorem $3.1$], where involutive solutions with primitive permutation group were completely classified, we focus on blocks of imprimitivity of involutive solutions (not necessarily simple). In addition to the main result of [15], which ensures the existence of a complete blocks system, we will show that a complete blocks system can be always obtained by the orbits of a minimal normal subgroup of the permutation group. To develop a theory useful to simple involutive solutions, in the second part we specialize to the irretractable ones. By tools that come from quasigroups theory, we decribe all the complete blocks system of irretractable involutive solutions with regular displacement group. Among these complete blocks systems, we distinguish the ones that give rise to congruences, providing an extension of a result involving the family of finite affine quasigroups (see [8, Lemma 4.8]) to a larger class of finite quasigroups. In order to understand more about congruences, using tools that come from the left quasigroups theory provided in [6], we develop a covering theory of left quasigroups that allows to recover the one given in [40] for indecomposable involutive solutions as a special case. This is the reason for which, throught the paper, we will use the language of cycle sets. Recall that a cycle set is a left quasigroup $(X,\cdotp)$ in which the equality $(x\cdotp y)\cdotp(x\cdotp z)=(y\cdotp x)\cdotp(y\cdotp z)$ holds for all $x,y,z\in X$. Cycle sets are useful to detect solutions of the Yang-Baxter equation: indeed, non-degenerate cycle sets bijectively correspond to non-degenerate involutive solutions, and by [33, Proposition 1-2] the correspondence is given by
@@
As one can expect, indecomposable (resp. simple) involutive solutions correspond to indecomposable (resp. simple) cycle sets. In the core of the paper, we focus on simple involutive solutions $(X,r)$ with nilpotent regular displacement group. We will show that, in this setting, $X$ must have a prime-power cardinality $p^{n}$ and the $p$-Sylow subgroup of the permutation group $\mathcal{G}(X)$ is normal. In this case we can reduce the classification to solutions for which $\mathcal{G}(X)$ is a $p$-group A complete description of these solutions will be provided, while the ones having cyclic displacement group will be classified. These results will be refined if, in addition, the displacement group is abelian and normal in the so-called total permutation group. In this last case, we exhibit a connection with irreducible representations of the First Weyl Algebra over a field with a prime number of elements. We remark that the First Weyl Algebra has already appeared in the study of involutive solutions, see [8, 38]. These solutions has always size $p^{k}$, for a prime number $p$ and a natural number $k$ divisible by $p$. Among these solutions, the ones having size $p^{p}$ will be completely classified. Finally, we illustrate our main results with some examples, and we use some of them to answer in the affirmative to [16, Question 7.6].
@@ 1.1 Basics
Note that $\mathcal{G}^{\alpha}$ is nothing but the kernel of the induced group homomorphism from $\mathcal{G}(X)$ onto $\mathcal{G}(X/\alpha)$ defined in Lemma 1 and it is admissible according to [6, Corollary 1.9]. On the other hand, given an admissible subgroup $H$, we can define the equivalence relation
@@
namely the equivalence relation with blocks given by the orbits of $H$. Such equivalence is a congruence [6, Lemma 1.8]. We usually denote the factor left quasigroup with respect to $\mathcal{O}_{H}$ just by $X/H$.
Moreover, given a congruence $\alpha$ we can define the subgroup $\mathrm{Dis}_{\alpha}$ as the normal closure of $\{\sigma_{x}\sigma_{y}^{-1}\,:\,x\,\alpha\,y\}$ in the group $\mathcal{G}(X)$. Such subgroups are admissible. We denote $\mathrm{Dis}_{1_{X}}$ just by $\mathrm{Dis}(X)$ and we refer to it as the displacement group of $X$. By [6, Section $1$], the quotient $\mathcal{G}(X)/\mathrm{Dis}(X)$ is cyclic, and in particular $\mathcal{G}(X)=\mathrm{Dis}(X)\langle\sigma_{x}\rangle$, for every $x\in X$.
@@ 1.2 Coverings of indecomposable left quasigroups
In this section we develop a covering theory for indecomposable left quasigroup which is consistent with the one developed in [37, Section 2] for indecomposable cycle sets.
By Lemma 1, the following definition makes sense. Clearly, it is consistent with [37, Definition 1].
@@
$Y\longrightarrow X,\quad g\mapsto g(x)$
@@
Following the terminology of [37, Section 2] and [40, Section 3], $u$ will be called universal covering. Now, we give the analog of [40, Proposition 4] for arbitrary indecomposable left quasigroups.
@@
The following theorem, which is the main result of the section, is an extension of [37, Theorem 1] to indecomposable left quasigroups and allows to understand the structure of the epimorphisms.
@@ 1.3 Blocks of imprimitivity
Let $G$ be a group acting on a set $X$ and $\alpha$ be an equivalence relation of $X$. We say that $\alpha$ is $G$-invariant if $g(x)\,\alpha\,g(y)$ provided $x\,\alpha\,y$. If $G$ is transitive, a subset $Y\subseteq X$ is called a block of imprimitivity if $g(Y)\cap Y$ is either empty or equal to $Y$, for every $g\in G$, and the partition $P$ of $X$ given by $P:=\{g(Y)\}_{g\in G}$ is said to be a complete blocks system. If the action of $G$ has no blocks of imprimitivity it is said to be primitive. If the action of $\mathcal{G}(X)$ is primitive we say that $X$ is primitive.
Let $X$ be a set and $H$ be a group of permutations on the set $Q$. We define
@@
Let $G$ be a group acting on a set $X$ and $H\trianglelefteq G$. Then $\mathcal{O}_{H}$ and $\mathcal{O}_{H}$ and $\sim_{H}$ are $G$-invariant equivalence relation on $Q$.
@@
    If $\sim_{H}=0_{Q}$ then $N_{\mathcal{G}(X)}(H_{x})\leq\mathcal{G}(X)_{x}$ for every $x\in X$. In particular $Z(\mathcal{G}(X))=1$.
@@
Let $X$ be a indecomposable left quasigroup and $H$ be a normal subgroup of $\mathcal{G}(X)$. Then $\sim_{H}=0_{Q}$ if and only if $N_{\mathcal{G}(X)}(H_{x})\leq\mathcal{G}(X)_{x}$ for every $x\in X$.
@@ 2 Left braces
Following [14, Definition 1], a set $B$ endowed of two operations $+$ and $\circ$ is said to be a left brace if $(B,+)$ is an abelian group, $(B,\circ)$ a group, and
@@
$\lambda_{x}(y):=-x+x\circ y,$
for all $y\in B$.
Let us recall the properties of the maps defined in (2).
@@
[34, Proposition 2],[14, Lemma 1] Let $B$ be a left brace. Then, the following are satisfied:
@@
For the following definition, we refer the reader to [34, pg. 160] and [14, Definition 3].
@@
The set $\{0\}$ is an ideal and it. will be called the trivial ideal. A left brace $B$ which contains no ideals different from $\{0\}$ and $B$ will be called a simple left brace.
A standard ideal of a left brace $B$, given in [34, Corollary of Proposition 6] and indicated by $B^{2}$, is the one given by the additive subgroup generated by the set $\{x*y\hskip 2.84526pt|\hskip 2.84526ptx,y\in B\}$, where $x*y:=-x+x\circ y-y$ for all $x,y\in B$.
In general, if $B$ is a left brace and $x\in B$ the map $\lambda_{x}$ does not belong to $Aut(B,+,\circ)$. The following lemma, that contains the same idea of [21, Lemma $1.1$], shows that under additional hypothesis, a map $\lambda_{x}$ induces an automorphism on some left ideals.
@@ 3.1 Basics
    If $X$ is a cycle set, then every deformation of $X$ is again a cycle set (see [39, Section $3$]).
@@
    A notion of displacement group for cycle sets was given in [8, Section $2$]. Even if it is slightly different from the one given in Section $1$ for arbitrary left quasigroup, by [8, Lemma $2.15$ and Proposition $2.17$] these two definitions coincide for non-degenerate cycle sets.
Let $B$ be a left brace. A cycle base is a subset $X$ of $B$ that is a union of orbits with respect to the $\lambda$-action and generating the additive group $(B,+)$. If a cycle base is a single orbit, then is called transitive cycle base. By [37, Theorem 3], every indecomposable cycle sets can be constructed by transitive cycle bases of left braces; on the other hand, every cycle set $X$ gives rise to a left brace structure on $\mathcal{G}(X)$ having as $\circ$ operation the usual maps composition (see Section $1$ of [37]).
@@
The following result, implicitly contained in [37], allows to avoid the construction of the previous result to recover indecomposable and irretractable cycle sets.
@@
By [40, Section 3], we know that if $X$ is an indecomposable cycle set and $I$ is an ideal of the left brace $\mathcal{G}(X)$, the $I$-orbits on $X$ give rise to a congruence of the cycle set $X$. This result can be recovered by looking at the connection between ideals and admissible subgroups.
@@
The statement follows by 20 and [6, Lemma 1.8]. ∎

### v2

@@ Introduction
The seminal papers due by Gateva-Ivanova and Van Den Bergh [26] and Etingov, Schedler, and Soloviev [25] encouraged several authors to the study of the involutive non-degenerate set-theoretic solutions (which we simply call involutive solutions). Recall that a solution $(X,r)$ is said to be decomposable if $X$ can be expressed as a disjoint union of two non-empty subsets $Y_{1}$ and $Y_{2}$ such that $r(Y_{i}\times Y_{i})\subseteq Y_{i}\times Y_{i}$ for all $i\in\{1,2\}$, and $(X,r)$ is called indecomposable, otherwise. To attack the classification-problem of involutive solutions, as first attempt one can restrict the investigations to the class of indecomposable ones. This approach is motivated by the fact that indecomposable solutions are in some sense the fundamental blocks that allow to understand all the involutive ones (see [25, Section $3$]). Even if we focus on indecomposable involutive solutions, a full classification is still hard, and in the last few years several authors investigated the so-called simple indecomposable involutive solutions, where a solution $(X,r)$ is said to be simple if every epimorphic image $(Y,s)$ is isomorphic to $(X,r)$ or is a singleton. A notion of simple involutive solution (not necessarily indecomposable) was at first introduced by Vendramin in [42] for involutive solutions not necessarily indecomposable. Even if his definition of simplicity is different, it coincides in the indecomposable case. Every simple involutive solution of size bigger than $2$ is indecomposable by [17]. These solutions play a special role since they are the “fundamental blocks” to construct all the indecomposable involutive ones by dynamical extensions (see [12, Proposition 2] and [42, Corollary 2.13]).
Even if the construction of simple involutive solutions appeared quite hard, several interesting results were obtained in the last few years. If $p$ is an odd prime number there exist a unique simple involutive solution of size $p$, see [25, Theorem 2.13]. For every natural number $n>1$, a simple involutive solution of cardinality $p^{n}$ was recently constructed in [19], and in the case $n=2$ a full classification was given in [22]. Further examples of cardinality $n^{2}$ and $mn^{2}$, where $n,m$ are arbitrary natural number, were constructed in [17, 20]. On the other hand, we know that an indecomposable involutive solutions of square-free cardinality can not be simple, as showed in [18].
@@
holds for all $x,y,z\in B$. The notion of simplicity was recently investigated in the general setting of bijective non-degenerate solutions (simply called solution) not necessarily involutive. In [13], a criterion of simplicity was provided by blocks of imprimitivity of the associated permutation group. A remarkable generalization of [10, Theorem $12$] was recently provided in [21], where the simple bijective solutions (not necessarily involutive) were completely characterised by the underlying permutation skew left braces, and several examples of non-involutive simple solutions were provided. In this way, the main Theorem of [21, Section $1$] represents a unifying tool to understand the notion of simplicity of several algebraic structures already present in literature. By these last results, it is clear that for the simplicity of a solution the ideal $B^{(2)}$ of the permutation left brace plays a fundamental role. However, we are still far away from a concrete description of the maps $r$ related to simple solutions. Indeed, in [21] a complete description of simple solutions provided by skew left braces $B$ with $B^{(2)}=0$ or $B^{(3)}=0$ was given, but the remaining case $B^{(2)}=B^{(3)}\neq 0$ is still open. As noted in [21, Section $1$], the last case includes the involutive simple solutions that are not of Lyuabashenko type. To partially fill this gap, one of the main goals of this paper is the study of finite involutive simple solutions in which $B^{(2)}$, that in this case coincides with the so-called displacement group, acts regularly on the underlying set. We will show that these solutions belong to the family of latin solutions, i.e. solutions in which the maps $\lambda_{x}$ provide a quasigroup structure on the set $X$ (see [9]).However, our investigation is not limited to simple solutions. Motivated by [16, Theorem $3.1$], where involutive solutions with primitive permutation group were completely classified, we focus on blocks of imprimitivity of involutive solutions (not necessarily simple). In addition to the main result of [16], which ensures the existence of a complete blocks system, we will show that a complete blocks system can be always obtained by the orbits of a minimal normal subgroup of the permutation group. To develop a theory useful to simple involutive solutions, in the second part we specialize to the irretractable ones. By tools that come from quasigroups theory, we decribe all the complete blocks system of irretractable involutive solutions with regular displacement group. Among these complete blocks systems, we distinguish the ones that give rise to congruences, providing an extension of a result involving the family of finite affine quasigroups (see [9, Lemma 4.8]) to a larger class of finite quasigroups. In order to understand more about congruences, using tools that come from the left quasigroups theory provided in [7], we develop a covering theory of left quasigroups that allows to recover the one given in [41] for indecomposable involutive solutions as a special case. This is the reason for which, throught the paper, we will use the language of cycle sets. Recall that a cycle set is a left quasigroup $(X,\cdotp)$ in which the equality $(x\cdotp y)\cdotp(x\cdotp z)=(y\cdotp x)\cdotp(y\cdotp z)$ holds for all $x,y,z\in X$. Cycle sets are useful to detect solutions of the Yang-Baxter equation: indeed, non-degenerate cycle sets bijectively correspond to non-degenerate involutive solutions, and by [34, Propositions 1-2] the correspondence is given by
@@
As one can expect, indecomposable (resp. simple) involutive solutions correspond to indecomposable (resp. simple) cycle sets. In the core of the paper, we focus on simple involutive solutions $(X,r)$ with nilpotent regular displacement group. We will show that, in this setting, $X$ must have a prime-power cardinality $p^{n}$ and the $p$-Sylow subgroup of the permutation group $\mathcal{G}(X)$ is normal. In this case we can reduce the classification to solutions for which $\mathcal{G}(X)$ is a $p$-group. A complete description of these solutions will be provided, while the ones having cyclic displacement group will be concretely classified. These results will be refined if, in addition, the displacement group is abelian and normal in the so-called total permutation group. In this last case, we exhibit a connection with irreducible representations of the First Weyl Algebra over a field with a prime number of elements. We remark that the First Weyl Algebra has already appeared in the study of involutive solutions, see [9, 39]. These solutions has always size $p^{k}$, for a prime number $p$ and a natural number $k$ divisible by $p$. Among these solutions, the ones having size $p^{p}$ will be completely classified. Finally, we illustrate our main results with some examples, and we use some of them to answer in the affirmative to [17, Question 7.6].
@@ 1.1 Basics
Note that $\mathcal{G}^{\alpha}$ is nothing but the kernel of the induced group homomorphism from $\mathcal{G}(X)$ onto $\mathcal{G}(X/\alpha)$ defined in Lemma 1 and it is admissible according to [7, Corollary 1.9]. On the other hand, given an admissible subgroup $H$, we can define the equivalence relation
@@
namely the equivalence relation with blocks given by the orbits of $H$. Such equivalence is a congruence [7, Lemma 1.8]. We usually denote the factor left quasigroup with respect to $\mathcal{O}_{H}$ just by $X/H$.
Moreover, given a congruence $\alpha$ we can define the subgroup $\mathrm{Dis}_{\alpha}$ as the normal closure of $\{\sigma_{x}\sigma_{y}^{-1}\,:\,x\,\alpha\,y\}$ in the group $\mathcal{G}(X)$. Such subgroups are admissible. We denote $\mathrm{Dis}_{1_{X}}$ just by $\mathrm{Dis}(X)$ and we refer to it as the displacement group of $X$. By [7, Section $1$], the quotient $\mathcal{G}(X)/\mathrm{Dis}(X)$ is cyclic, and in particular $\mathcal{G}(X)=\mathrm{Dis}(X)\langle\sigma_{x}\rangle$, for every $x\in X$.
@@ 1.2 Coverings of indecomposable left quasigroups
In this section we develop a covering theory for indecomposable left quasigroup which is consistent with the one developed in [38, Section 2] for indecomposable cycle sets.
By Lemma 1, the following definition makes sense. Clearly, it is consistent with [38, Definition 1].
@@
$u:Y\longrightarrow X,\quad g\mapsto g(x)$
@@
Following the terminology of [38, Section 2] and [41, Section 3], $u$ will be called universal covering. Now, we give the analog of [41, Proposition 4] for arbitrary indecomposable left quasigroups.
@@
The following theorem, which is the main result of the section, is an extension of [38, Theorem 1] to indecomposable left quasigroups and allows to understand the structure of the epimorphisms.
@@ 1.3 Blocks of imprimitivity
Let $G$ be a group acting on a set $X$ and $\alpha$ be an equivalence relation of $X$. We say that $\alpha$ is $G$-invariant if $g(x)\,\alpha\,g(y)$ provided $x\,\alpha\,y$. If $G$ is transitive, a subset $Y\subseteq X$ is called a block of imprimitivity if $g(Y)\cap Y$ is either empty or equal to $Y$, for every $g\in G$, and the partition $P$ of $X$ given by $P:=\{g(Y)\}_{g\in G}$ is said to be a complete blocks system. If the action of $G$ has no non-trivial blocks of imprimitivity it is said to be primitive. If the action of $\mathcal{G}(X)$ is primitive we say that $X$ is primitive.
Let $X$ be a set and $H$ be a group of permutations on the set $X$. We define
@@
Let $G$ be a group acting on a set $X$ and $H\trianglelefteq G$. Then $\mathcal{O}_{H}$ and $\mathcal{O}_{H}$ and $\sim_{H}$ are $G$-invariant equivalence relation on $X$.
@@
    If $\sim_{H}=0_{X}$ then $N_{\mathcal{G}(X)}(H_{x})\leq\mathcal{G}(X)_{x}$ for every $x\in X$. In particular $Z(\mathcal{G}(X))=1$.
@@
Let $X$ be a indecomposable left quasigroup and $H$ be a normal subgroup of $\mathcal{G}(X)$. Then $\sim_{H}=0_{X}$ if and only if $N_{\mathcal{G}(X)}(H_{x})\leq\mathcal{G}(X)_{x}$ for every $x\in X$.
@@ 2 Left braces
Following [15, Definition 1], a set $B$ endowed of two operations $+$ and $\circ$ is said to be a left brace if $(B,+)$ is an abelian group, $(B,\circ)$ a group, and
@@
$\lambda_{x}(y):=-x+x\circ y,$ (5)
for all $y\in B$. Let us recall the properties of the maps $\lambda_{x}$.
@@
[35, Proposition 2],[15, Lemma 1] Let $B$ be a left brace. Then, the following are satisfied:
@@
For the following definition, we refer the reader to [35, pg. 160] and [15, Definition 3].
@@
The set $\{0\}$ is an ideal and it will be called the trivial ideal. A left brace $B$ which contains no ideals different from $\{0\}$ and $B$ will be called a simple left brace.
A standard ideal of a left brace $B$ is the socle, indicated by $Soc(B)$ and given by the kernel of the map $\lambda$. Another useful ideal of a left brace $B$, given in [35, Corollary of Proposition 6] and indicated by $B^{2}$, is the one given by the additive subgroup generated by the set $\{x*y\hskip 2.84526pt|\hskip 2.84526ptx,y\in B\}$, where $x*y:=-x+x\circ y-y$ for all $x,y\in B$. In general, if $I$ is an ideal of a left brace $B$, then $I*B$ and $B*I$ are subset of $I$, where $I*B$ and $B*I$ are defined in the same way of $B^{2}$.
In general, if $B$ is a left brace and $x\in B$ the map $\lambda_{x}$ does not belong to $Aut(B,+,\circ)$. The following lemma, that contains the same idea of [22, Lemma $1.1$], shows that under additional hypothesis, a map $\lambda_{x}$ induces an automorphism on some left ideals.
@@ 3.1 Basics
    If $X$ is a cycle set, then every deformation of $X$ is again a cycle set (see [40, Section $3$]).
@@
    A notion of displacement group for cycle sets was given in [9, Section $2$]. Even if it is slightly different from the one given in Section $1$ for arbitrary left quasigroup, by [9, Lemma $2.15$ and Proposition $2.17$] these two definitions coincide for non-degenerate cycle sets.
Let $B$ be a left brace. A cycle base is a subset $X$ of $B$ that is a union of orbits with respect to the $\lambda$-action and generating the additive group $(B,+)$. If a cycle base is a single orbit, then is called transitive cycle base. By [38, Theorem 3], every indecomposable cycle sets can be constructed by transitive cycle bases of left braces; on the other hand, every cycle set $X$ gives rise to a left brace structure on $\mathcal{G}(X)$ having as $\circ$ operation the usual maps composition (see Section $1$ of [38]).
@@
The following result, implicitly contained in [38], allows to avoid the construction of the previous proposition to recover indecomposable and irretractable cycle sets.
@@
By [41, Section 3], we know that if $X$ is an indecomposable cycle set and $I$ is an ideal of the left brace $\mathcal{G}(X)$, the $I$-orbits on $X$ give rise to a congruence of the cycle set $X$. This result can be recovered by looking at the connection between ideals and admissible subgroups.
@@
The statement follows by 20 and [7, Lemma 1.8]. ∎